Hello Tutor can you please help me with this question the product of 2 consecutive odd integers is equal to 30 more than the first.Find the integors
Found 2 solutions by lynnlo, ikleyn:
Answer by lynnlo(4176)
Answer by ikleyn(53748)

.
the product of 2 consecutive odd integers is equal to 30 more than the first. Find the integers
~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Let your lesser odd number be 'n'.
Then the other odd number is (n+2).


Your equation is

    n*(n+2) = n + 30.


Simplify

    n^2 + 2n = n + 30,

    n^2 + n - 30 = 0,

    (n+6)*(n-5) = 0.


The roots are -6 and 5.


You want the odd number - so, you select 5 and reject -6.


ANSWER.  The numbers are 5 and 7.

Solved.

--------------------

Ignore the answer by @lynnlo, since it is incorrect.